OpenStudy (anonymous):

How do I convert this quadratic equation into vertex form? It was originally in intercept form as f(x) = (x + 6)(x - 2), but I now have it in standard form as f(x) = x^2 + 4x - 12. I just need help converting to vertex form.

OpenStudy (anonymous):

I know it has something to do with "completing the square" but I'm not quite sure how to do that.

hartnn (hartnn):

what is the co-efficient of x ? (we are starting completing the square)

OpenStudy (anonymous):

4... right?

hartnn (hartnn):

correct! divide it by 2 and square it, what u get ?

OpenStudy (anonymous):

4/2 = 2, and 2 squared is 4 again

hartnn (hartnn):

right , so add and subtract 4

hartnn (hartnn):

(x^2+4x+4)-4-12 got what i did?

hartnn (hartnn):

my final aim is to bring it in vertex form \(\Large f(x)=a(x-h)^2+k \)

OpenStudy (anonymous):

Oh yeah I see now, took me a minute lol

OpenStudy (anonymous):

So (x^2 + 4x + 4) - 4 - 12 is simplified into (x^2 + 4x + 4) - 16... and (x^2 + 4x + 4) can be factored into (x + 2)^2?

OpenStudy (anonymous):

I was told vertex form is y - k = a(x - h)^2

hartnn (hartnn):

same thing. and yes f(x)=(x+2)^2-16 is correct vertex form :)

OpenStudy (anonymous):

okay thank you (:

OpenStudy (anonymous):

so the x-value of the vertex would be -2?

hartnn (hartnn):

yup.

hartnn (hartnn):

welcome ^_^